OpenStudy (anonymous):

a spinner and number cube are used in a game.The spinner has an equal chance of landing on one of five colors: red,yellow,blue,green and purple. the faces of number cube are labeled 1to 6. 1. What is the probability of spinning blue then rolling 3 or 4? 2. What is the probability of rolling an even number and spinning yellow? 3. What is the probability of a rolling a 2 or 4 on the number cube and not landing on red on the spinner?

OpenStudy (anonymous):

1. 1/ 5 * 2/6 = 1/15 2. 1/5 * 3/6 = 1/10 3. 4/5 * 2/6 = 4/15

OpenStudy (anonymous):

is that all three questions answer

OpenStudy (anonymous):

yes.......

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

you welcome..........

OpenStudy (anonymous):

i brb

OpenStudy (anonymous):

for getting explanation you have to talk ................

OpenStudy (anonymous):

younbird r u there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

did u understand it...???